OpenStudy (anonymous):

could someone help me with this problem? -2(x+4)+8(3x-1)

OpenStudy (anonymous):

could u also show your work?

OpenStudy (saifoo.khan):

\[-2(x+4) + 8(3x-1)\]Using distributive prop, \[-2x - 8 + 24x - 8\]Add like terms, \[22x - 16\]

OpenStudy (edchau3):

−2(x+4)+8(3x−1) Using distributive prop, −2x−8+24x−8 Add like terms, 22x−16 Divide both sides by 22, 22x/22=16/22 Solve for x x=0.7272727272727273